Heroux Corporation has two manufacturing departments--Forming and Customizing. The company used the following data at the beginning of the year to calculate predetermined overhead rates: Forming Customizing Total Estimated total machine-hours (MHs) 3,000 7,000 10,000 Estimated total fixed manufacturing overhead cost $ 16,500 $ 20,300 $ 36,800 Estimated variable manufacturing overhead cost per MH $ 1.70 $ 2.50 During the most recent month, the company started and completed two jobs--Job A and Job H. There were no beginning inventories. Data concerning those two jobs follow: Job A Job H Direct materials $ 12,800 $ 6,700 Direct labor cost $ 24,300 $ 7,800 Forming machine-hours 2,000 1,000 Customizing machine-hours 2,800 4,200 Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. The manufacturing overhead applied to Job H is closest to:

  • $29,880

  • $22,680

  • $30,888

  • $7,200
Answer #1

Overhead rate for forming = (16500/3000+1.7) = 7.2

Overhead rate for customizing = (20300/7000+2.5) = 5.4 per hour

Overhead applied = (1000*7.2+4200*5.4) = 29880

So answer is a) $29880